OpenStudy (anonymous):

A scientist is doing an experiment with a chemical that has a half-life of 11 days. How much of the chemical will be remaining after 40 days if there were 300 grams at the beginning of his lab?

OpenStudy (lgbasallote):

the formula for half life is \[\frac{\ln 2}{k} = t\] so \[k = \frac{\ln 2}{t}\] \[\implies k = \frac{\ln 2}{11}\] \[\implies k = 0.0630\] to find how much there would be after 40 days use the exponential decay formula \[x = x_o e^{-kt}\] \[\implies x = (300)(e)^{-0.0630 \times 40}\] \[\implies x = (300)e^{-2.52}\] \[\implies x = (300)(0.0805)\] \[\implies x = 24.14\] \[\implies x \approx 25\] does that help?

OpenStudy (anonymous):

yes thanks

OpenStudy (lgbasallote):

welcome